Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

PROBABILITY QUESTION (PLEASE HELP!!!)

Expert replies
by Stockmoose16 » Tue Sep 09, 2008 2:38 pm
I desperately need someone to explain the difference between these two questions:

In a bag of 10 marbles, there are 6 white, 4 blue:

1) "What is the probability of getting a white marble and blue marble"

2) "What is the probability of first picking a white marble, then picking a blue one?

These two questions sound exactly the same to me, but according to my GMAT book, have two different answers. Can someone explain the difference between the two questions in layman's terms?
Join the discussion
Source: — Problem Solving |

Stockmoose16 wrote:I desperately need someone to explain the difference between these two questions:

In a bag of 10 marbles, there are 6 white, 4 blue:

1) "What is the probability of getting a white marble and blue marble"

2) "What is the probability of first picking a white marble, then picking a blue one?

These two questions sound exactly the same to me, but according to my GMAT book, have two different answers. Can someone explain the difference between the two questions in layman's terms?
"What is the probability of getting a white marble and blue marble"

= 6/10 * 4/9 + 4/10 * 6/9

= 4/15 + 4/15 = 8/15 [here order of picking marbles is not mentioned, therefore you have consider both the possibilities]

"What is the probability of first picking a white marble, then picking a blue one?

= 6/10*4/9 = 4/15


Hope this helps.
Join the discussion

parallel_chase wrote:
Stockmoose16 wrote:I desperately need someone to explain the difference between these two questions:

In a bag of 10 marbles, there are 6 white, 4 blue:

1) "What is the probability of getting a white marble and blue marble"

2) "What is the probability of first picking a white marble, then picking a blue one?

These two questions sound exactly the same to me, but according to my GMAT book, have two different answers. Can someone explain the difference between the two questions in layman's terms?
"What is the probability of getting a white marble and blue marble"

= 6/10 * 4/9 + 4/10 * 6/9

= 4/15 + 4/15 = 8/15 [here order of picking marbles is not mentioned, therefore you have consider both the possibilities]

"What is the probability of first picking a white marble, then picking a blue one?

= 6/10*4/9 = 4/15


Hope this helps.
Can you please explain why the order makes any difference? (effectively doubling the answer?)
Join the discussion

Stockmoose16 wrote: Can you please explain why the order makes any difference? (effectively doubling the answer?)
The order makes a difference in probability since we are not given any order of picking marbles, like there was in second part of the question.

We are not doubling anything. We are just calculating the total probable combination of picking marbles.
Join the discussion